Consider a 6 story building with a total plan area B x L. Which B=100ft and L=60ft. The average weight of each floor and roof are 100lb/ft2 and 80lb/ft2, respectively. The 72ft tall building is located in a seismic region with SDS= 1.0g and SD1= 0.7 for a building supported on rock, where g is the gravitational acceleration. The structure of the building consists of steel moment frames (all joints are rigid) that have R value of 8. The importance factor 1.1. Determine the total design seismic force (kips) acting at the base of the building (seismic base shear in kips).
Answer #1

Solution:-

Given

L= 60 feet

B = 100 feet

Floor weight = 100 lb/feet2

= 100*100*60 = 600 kips

Weight of roof = 80 lb/feet2

= 80*100*60 = 480 kips

Total weight of building(W) = 5*600 + 480 = 3480 kips

Design base shear (VB) = Ah*W

Horizontal seismic coefficient (Ah) = Z/2 * I/R *Sa/g

= 0.24/2*1.1/8 *1 = 0.0165

Total seismic design base shear(VB) = Ah*W

= 0.0165 * 3480

VB = 57.42 kips   Answer